OpenStudy (anonymous):

Leo has a salary of $120,000. Calculate his new salary if it is : (a) increased by 10%, then decreased by 10% (b) increased by x%, then decreased by x% Please help!! Thanks!~

OpenStudy (kropot72):

(a) New salary is $120,000 * 1.1 * 0.9 = you can calculate

hartnn (hartnn):

hi :) did u get that ? or need some more explanation ?

OpenStudy (kropot72):

The solution to (a) can be put in more detail as follows: \[New\ salary=$120,000\times(1+\frac{10}{100})\times(1-\frac{10}{100})\]

OpenStudy (anonymous):

I got part (a) --Thanks @kropot72--, but I don't know how to get the answer for part (b). According to my textbook, the answer for part (b) is $120,000 ( 1- x^2/10^4) Please explain!!

OpenStudy (kropot72):

(b) If you look at the detailed posting of the solution to (a) \[New\ salary=$120,000\times(1+\frac{10}{100})\times(1-\frac{10}{100})\] and plug in x instead of 10 we get \[New\ salary=$120,000\times(1+\frac{x}{100})(1-\frac{x}{100})\] Can you see that the two terms in x are the factorization of the difference of two squares?

OpenStudy (anonymous):

Yes~

OpenStudy (kropot72):

So when we multiply out those two terms in x we get \[1-(\frac{x}{100})^{2}\]

hartnn (hartnn):

lets take it part by part for any quantity say, A, 10% of it is A*10/100 x% of A is A*x/100 If A increased by 10% means A + 10% of A = A (1+ 10/100) If A increased by x% means A + x% of A = A (1+ x/100) thats the basic thing, you can apply this repeatedly in your problem

OpenStudy (kropot72):

\[1-(\frac{x}{100})^{2}=(1-\frac{x ^{2}}{10^{4}})\]

hartnn (hartnn):

for decrease, you just use negative sign, A decreased by x% = A (1-x/100)

OpenStudy (anonymous):

Oh, I get it now! Thanks a lot @kropot72 and @hartnn :)

OpenStudy (kropot72):

You're welcome :)